OpenStudy (anonymous):

4^3square root of 10 + 2 ^3 square root of 10 can some one please help me

OpenStudy (zehanz):

You probably do not mean: \[4^3\sqrt{10}+2^3\sqrt{10}\]although this is what you wrote. Could you be more specific?

OpenStudy (zehanz):

If this IS what you meant, however, then because 4³=4x4x4=64 and 2³=2x2x2=8, you've got a total of \[(64+8)\cdot \sqrt{10}=72\sqrt{10}\]